OpenStudy (anonymous):

What is the volume of a cone with a radius of 3 and a height of 19? A. 1083 units3 B. 537.21 units3 C. 39.79 units3 D. 179.07 units

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } \pi r ^{2} h\]

OpenStudy (anonymous):

57 pi

OpenStudy (anonymous):

I Dont Kno How To Do It Then lo Its Getting hARD iM In Geometry

OpenStudy (anonymous):

Ok, so plug in 3 for r. You should get (3)^2. That = 9.

OpenStudy (anonymous):

Your h is 19.

OpenStudy (anonymous):

So now, you just plug all them in. (1/3)(pi)(9)(19)

OpenStudy (anonymous):

Does she want you to use 3.14 or 22/7?

OpenStudy (anonymous):

3.14

OpenStudy (anonymous):

So what did you get?

OpenStudy (anonymous):

I Dont Have A Fraction Button On my Calculator

OpenStudy (anonymous):

What Is The Decimal For 1/3

OpenStudy (anonymous):

.33333333333333333 and on and on

OpenStudy (anonymous):

Thanks It Is d

OpenStudy (anonymous):

Can U Keep Helping Me? Cause I Only hAVE 5 Quizes Left

OpenStudy (anonymous):

I Graduate Today

OpenStudy (anonymous):

Which of the following is a solid bounded by the set of all points at a given distance from a given point? A. Cube B. Sphere C. Cone D. Cylinder

OpenStudy (anonymous):

I Graduate Today

OpenStudy (anonymous):

It should be whatever is a constant radius from the center...

OpenStudy (anonymous):

Thanks Its B

OpenStudy (anonymous):

Tiana would like to buy cone-shaped hats for her birthday party. She wants hats with the largest volume, since they look bigger. Which of the following brands of hat should she buy? A. Brand A B. Brand B C. Either — the hats have equal volumes

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Are there pictures?

OpenStudy (anonymous):

I Took A Picture Of It

OpenStudy (anonymous):

Find the volume of each :) The same formula I gave you a little while ago, use that for each one. The one with the largest value is the answer.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The ratio of the lengths of corresponding parts in two similar solids is 5:1. What is the ratio of their surface areas? A. 125:1 B. 10:1 C. 5:1 D. 25:

OpenStudy (anonymous):

It Is d But What Is This Which expression gives the volume of a sphere with radius 17? A. (173) B. (172) C. 4(172) D. 4(173)

OpenStudy (anonymous):

The formula for volume of a sphere is (4/2)(pi)(r^3). So r^3, what is 17^3?

OpenStudy (anonymous):

Ok I Got It Thanks Its c

OpenStudy (anonymous):

:) Good luck

OpenStudy (anonymous):

Thanks